OpenStudy (anonymous):

Ignacio earns 5 percent simple interest annually on his savings account. How much interest will he earn each month on $1250?

OpenStudy (anonymous):

I don't know if I undestood the question (my english is bad haha), but I think the answer is: =1250*0.05=62.5 Each month: 62.5/12=5.21